# Towards Full-scene Domain Generalization in Multi-agent Collaborative Bird’s Eye View Segmentation for Connected and Autonomous Driving

### v2

@@ Abstract
Collaborative perception has recently gained significant attention in autonomous driving, improving perception quality by enabling the exchange of additional information among vehicles. However, deploying collaborative perception systems can lead to domain shifts due to diverse environmental conditions and data heterogeneity among connected and autonomous vehicles (CAVs). To address these challenges, we propose a unified domain generalization framework applicable in both training and inference stages of collaborative perception. In the training phase, we introduce an Amplitude Augmentation (AmpAug) method to augment low-frequency image variations, broadening the model’s ability to learn across various domains. We also employ a meta-consistency training scheme to simulate domain shifts, optimizing the model with a carefully designed consistency loss to encourage domain-invariant representations. In the inference phase, we introduce an intra-system domain alignment mechanism to reduce or potentially eliminate the domain discrepancy among CAVs prior to inference. Comprehensive experiments substantiate the effectiveness of our method in comparison with the existing state-of-the-art works. Code will be released at https://github.com/DG-CAVs/DG-CoPerception.git.
@@ I Introduction
Recently, multi-agent collaborative perception liuWhen2comMultiAgentPerception2020 ; huWhere2commCommunicationEfficientCollaborative2022 ; yangWhat2commCommunicationefficientCollaborative2023 ; liuWho2comCollaborativePerception2020 ; xuV2XViTVehicletoEverythingCooperative2022 ; chenVehicleServiceVaaS2023 has attracted increasing attention in the autonomous driving community, due to its promising way to overcome the limitations of a single-agent perception system. For example, a single-agent perception system may suffer from occlusion and sparse sensor observation at a far distance, which may lead to a high risk of traffic accidents. Collaborative perception can address these issues by employing vehicle-to-vehicle (V2V) collaboration, where visual information (e.g., raw sensory information, intermediate perception features, and perception results) from multiple nearby CAVs can be shared to improve the accurate understanding of the environment.
@@
In collaborative perception, current approaches mainly aim to strike a balance between performance and bandwidth consumption while developing new perception architectures. For example, Liu et al. liuWho2comCollaborativePerception2020 leveraged a three-step handshake communication protocol to determine the information of which CAVs should be shared. Chen et al. chenCooperCooperativePerception2019 conducted a study to implement collaborative perception by fusing raw LiDAR point clouds of different CAVs. Xu et al. xuV2XViTVehicletoEverythingCooperative2022 proposed a new vision transformer framework which can fuse information of the on-road vehicles and the roadside infrastructure. Hu et al. senkangAdaptiveCommunicationsCollaborative2023 proposed an adaptive communication scheme which can construct the communication graph and minimize the communication delay according to different channel information states.
@@
While the previously mentioned methods have made significant progresses in balancing performance and bandwidth while optimizing communication graph construction, they regrettably overlook a critical challenge - the domain shift problem in collaborative perception. These methods typically assume that both training and testing data originate from the same domain, a condition rarely met in real-world applications. The consequence of neglecting the domain shift is a severe deterioration in perception performance when these methods are implemented in environments characterized by domain variations. For instance, in OPV2V xuOPV2VOpenBenchmark2022 and nuScenes caesarNuScenesMultimodalDataset2020 , the widely recognized benchmarks for bird’s eye view (BEV) semantic segmentation, their data collection settings exhibit notable disparities, encompassing differences in intrinsic and extrinsic parameters. One dataset is sourced from a simulator, while the other derives from real-world observations. When a model initially trained on OPV2V is directly transferred to the nuScenes dataset, a substantial performance drop is observed. Consequently, there is an urgent need to achieve domain generalization for collaborative perception.
@@
Domain generalization (DG) is an active research topic with numerous methods liDomainAdaptationBased2023 ; zhouDomainGeneralizationMixStyle2021 ; liDomainGeneralizationAdversarial2018 ; yangFDAFourierDomain2020 ; volpiGeneralizingUnseenDomains2018 . However, when applied to collaborative perception in the context of CAVs, the domain generalization problem becomes notably more intricate. Several challenges make collaborative perception in CAVs a demanding domain for DG. First, for collaborative BEV segmentation, there are few open-source benchmarks, and annotating data is resource-intensive and time-consuming. Only OPV2V xuOPV2VOpenBenchmark2022 and V2X-Sim liV2XSimMultiAgentCollaborative2022 provide accessible datasets. Moreover, CAV environment is highly dynamic, subject to frequent changes. For example, domains shift when transitioning from city roads to countryside, and similar shifts are also observed when moving from daytime to nighttime conditions. Furthermore, data heterogeneity among collaborative vehicles also presents a significant challenge. In practical scenarios, different vehicles may operate under diverse environmental conditions during collaborative perception. For instance, one vehicle might navigate in darkness, while another is positioned in a well-lit area. The absence of standardized parameters across different onboard cameras lead to variations in image quality, affecting factors like brightness, contrast, and color zhouDomainGeneralizationSurvey2022 . As a result, these divergent environmental conditions and sensor characteristics inevitably create a substantial domain gap between vehicles, leading to performance degradation in collaborative perception.
@@
To tackle these challenges, we present a comprehensive domain generalization framework that operates seamlessly in both the training and inference phases of collaborative perception when machine learning is used. Firstly, our approach starts with the development of an amplitude augmentation method, a resource-efficient technique that does not need extra annotated data and is ready for immediate use. Next, we introduce a meta-consistency training scheme rooted in meta-learning principle. This scheme equips the model with the ability to generalize effectively to unseen domains while fostering the acquisition of domain-invariant features. Finally, our framework includes an intra-system domain alignment scheme, strategically designed to mitigate data heterogeneity and bridge the domain gaps among various CAVs. The main contributions of this paper are summarized as follows.
@@
  • We address the practical problem of domain generalization for collaborative perception. To the best of our knowledge, this is the first work to study domain generalization in the area of collaborative perception.
@@
The remainder of this paper is organized as follows. In Sec. II, we review the related work of collaborative perception and domain generalization. In Sec. III, we introduce our proposed domain generalization method in detail. In Sec. IV, we conduct comprehensive experiments to evaluate our method. Finally, we conclude our work in Sec. V.
@@ II Related Work
II-A Collaborative Perception
Despite the notable progress in autonomous driving in recent years, single-agent perception systems face significant challenges related to occlusions and limitations of sensor coverage. For example, in a crowded urban environment, a vehicle may be blocked by other vehicles, pedestrians, or buildings, which makes it difficult to perceive its surrounding environment. Multi-agent collaborative perception emerged as a solution to addressing these issues wangV2VNetVehicletoVehicleCommunication2020 ; liuWho2comCollaborativePerception2020 ; liuWhen2comMultiAgentPerception2020 ; huWhere2commCommunicationEfficientCollaborative2022 . Based on data sharing strategies, collaborative perception can generally be divided into three aspects: 1) early fusion chenCooperCooperativePerception2019 where raw data is shared with ego vehicle; 2) intermediate fusion where intermediate features extracted from each vehicle’s sensor data are transmitted and shared; and 3) late fusion where perception outputs are shared (e.g., bounding box position, segmentation mask, etc.). Recent works indicate that intermediate fusion is more efficient than early fusion and late fusion because it effectively balances accuracy and communication bandwidth. For example, Liu et al. liuWho2comCollaborativePerception2020 proposed a handshake communication mechanism, which consists of three stages, namely request, match, and connect, which can determine which agents should communicate. Li et al. liLearningDistilledCollaboration2021 leveraged knowledge distillation to train a model by a teacher-student framework and proposed a kind of matrix that allows the agent to adaptively highlight the information region, thereby achieving effective intermediate fusion. Wang et al. wangV2VNetVehicletoVehicleCommunication2020 proposed an intermediate fusion strategy where all agents transmit features derived from the raw point cloud to strike the balance between bandwidth and precision. Despite the aforementioned advances of fusion methods, the domain shift caused by different environmental conditions and data heterogeneity among CAVs has not been investigated well in collaborative perception.
@@ II-B Domain Generalization
The domain shift problem has seriously hindered large-scale deployments of machine learning models. To tackle this issue, a lot of DG methods have been proposed gongDLOWDomainFlow2019 ; ghifaryDomainGeneralizationObject2015 ; liEpisodicTrainingDomain2019 ; yangFDAFourierDomain2020 ; liFeatureCriticNetworksHeterogeneous2019 . DG is dedicated to learning a model trained on multiple source domains that can generalize to unseen target domains. For example, Motiian et al. motiianUnifiedDeepSupervised2017 leveraged a contrastive loss to learn an embedding subspace that is discriminative, and the mapped visual domains are semantically aligned and yet maximally separated, which requires only a few labeled target samples. Some other domain generalization methods are based on data augmentation, which is a common practice to regularize the training of machine learning models to avoid overfitting and improve generalization, e.g., Volpi et al. volpiAddressingModelVulnerability2019 defined a new data augmentation rule to transform images with a different distribution where the model is most vulnerable, thereby making the model more robust against distribution shifts. Volpi et al. volpiGeneralizingUnseenDomains2018 also proposed a data augmentation method to leverage adversarial gradients obtained from a task classifier to perturb input images. In addition, meta-learning, also known as learn-to-learn, is another method helpful for domain generalization, which exposes a model to domain shift and guides the model to learn how to learn from the domain shift. Finn et al. finnModelAgnosticMetaLearningFast2017 proposed MAML, a method that partitions the training data into meta-train and meta-test subsets, and trains a model using the meta-train subset to enhance its performance on the meta-test subset. Our method also utilizes meta-learning scheme as our foundational learning scheme. However, we propose a novel amplitude augmentation method to simulate the domain shift, and leverage a consistency optimization to enable the model to learn domain-invariant features.
Figure 1: Overall Architecture. 1) The first part of our proposed method is amplitude augmentation which can transform the original source domain data to different target domains. 2) The second part is meta-consistency training, which can simulate the domain shift, guiding the model to learn how to learn from different domains. Then, we exploit the meta-consistency loss to encourage our model to learn the domain-invariant features, thereby enhancing the ability of generalization of the model. 3) The third part is intra-system domain alignment, which can minimize the domain gap among the data perceived by different collaborative vehicles prior to inference.
@@ III-A Problem Formulation and Overview
Collaborative Domain Generalization. Let $(\mathcal{X},\mathcal{Y})$ denote the input and label space in a collaborative BEV segmentation task, $\mathcal{D}=\{\mathcal{D}_{1},\mathcal{D}_{2},\cdots,\mathcal{D}_{K}\}$ be the set of $K$ source domains. In each domain, $\mathcal{D}_{k}=\{(x_{i}^{(k)},y_{i}^{(k)})\}^{N_{k}}_{n=1}$ is a set of samples from domain distribution ($\mathcal{X}_{k},\mathcal{Y}$), where $x_{i}^{(k)}$ is the input image set of CAVs at a certain time, $y_{i}^{(k)}$ is the corresponding label, $N_{k}$ is the number of samples in domain $k$. The goal is to learn a collaborative BEV segmentation model $F:\mathcal{X}\rightarrow\mathcal{Y}$ based on the data from multiple source domains $\mathcal{D}$, which can be generalized to the unseen target domain.
@@
Following AmpAug, we propose our meta-consistency training scheme, a method capable of simulating domain shifts and optimizing the model using a carefully crafted consistency loss function. This loss drives the encoder to learn domain-invariant features by minimizing the maximum mean discrepancy of the features obtained during both meta-training and meta-testing phases. Finally, our intra-system domain alignment mechanism is introduced, aimed at mitigating, or even eliminating, domain discrepancies among connected and autonomous vehicles (CAVs) before the inference stage. This is achieved by translating image styles and unifying the distribution of image pixels in color space. Each of these components will be discussed in detail in the following sections. The overall framework is shown in Fig. 1.
@@ III-B Amplitude Augmentation
Figure 2: Visualization of our constructed target dataset $\mathcal{D}_{t}$.
@@
As for collaborative perception, the environments sensed by CAVs are extraordinarily complex due to dynamic transformations across time and space. Factors such as varying weather conditions, diverse driving scenarios, fluctuating illuminations, and distinct image colors contribute to this complexity. These environmental fluctuations result in domain discrepancy, significantly impeding perception performance when deploying models in real-world settings.
@@
Our research, upon careful examination, has revealed that this domain discrepancy primarily arises from variations in the low-frequency spectrum of images. Importantly, these variations do not affect the image’s finer details, encompassing high-frequency spectrum elements like objects, cars, and people. Furthermore, these low-frequency spectrum variations hold substantial sway over the image. If these variances are not adequately represented in the training dataset, the models fail to generalize effectively. Consequently, we introduce a straightforward yet highly effective method known as AmpAug. AmpAug serves the purpose of augmenting these low-frequency variations, compelling the model to acquire domain-invariant representations. This augmentation, in turn, reduces the model’s susceptibility to the influence of low-frequency spectrum variations in images.
@@
First, we construct a dataset $\mathcal{D}_{t}$, which is carefully collected under different weather conditions (including sunny, rainy, foggy days), driving scenarios, colors, etc. The visualization of our constructed target dataset $\mathcal{D}_{t}$ is shown in Fig. 2. The reason for constructing this dataset is that we want to obtain enough different amplitude spectrum of the images and leverage it to augment the amplitude spectrum of the images in the source domain dataset $\mathcal{D}_{s}$.
@@
Specifically, consider our constructed dataset $\mathcal{D}_{t}=\{x_{i}^{t}\}^{N_{t}}_{i=1}$ and a source domain dataset $\mathcal{D}_{s}=\{x_{i}^{s},y_{i}^{s}\}^{N_{s}}_{i=1}$, where $x_{i}^{t},x_{i}^{s}\in\mathbb{R}^{H\times W\times C}$ ($H$ and $W$ are the height and width of the image, respectively; $C$ is the number of channels), $y_{i}^{s}\in\mathbb{R}^{H\times W\times C}$ is the corresponding label. We first sample an image $x_{i}^{t}$ from our constructed dataset $\mathcal{D}_{t}$, and convert it to the frequency space by fast Fourier transform (FFT):
$\mathbf{F}(x_{i}^{t})(u,v,c)=\sum^{H-1}_{j=0}\sum^{W-1}_{k=0}x_{i}^{t}(h,w,c)e^{-2j\pi\left(\frac{h}{H}u+\frac{w}{W}v\right)}$
@@
After obtaining the frequency representation $\mathbf{F}(x_{i}^{t})$, we can decompose it into amplitude spectrum $A(x_{i}^{t})\in\mathbb{R}^{H\times W\times C}$ and phase spectrum $P(x_{i}^{t})\in\mathbb{R}^{H\times W\times C}$, which represents the low-level distributions (e.g., style, light, etc.) and high-level distributions (e.g., details and object, etc.), respectively yangFDAFourierDomain2020 . Then, we sample another image $x_{j}^{s}$ from the source dataset and perform the same operation to obtain the amplitude spectrum $A(x_{j}^{s})$ and phase spectrum $P(x_{j}^{s})$. After that, we can obtain the augmented image $x_{j}^{s\rightarrow t}$ by incorporating the low-level distribution $A(x_{j}^{s})$ and $A(x_{i}^{t})$.
@@
We first introduce a binary mask $M$ which can be used to control the proportion of the low-level distribution $A(x_{j}^{s})$ and $A(x_{i}^{t})$ in the augmented image $x_{j}^{s\rightarrow t}$, $M=(h,w)=\mathbf{1}_{h\times w}$ where $h\in[-\alpha H:\alpha H]$, $w\in[-\alpha W:\alpha W]$, $\alpha\in(0,1)$. Then, we generate a new amplitude spectrum distribution by as follows:
$A(x_{j}^{s\rightarrow t})=(I-M)\cdot A(x^{s}_{j})+M\cdot A(x^{t}_{i})$ (1)

### v3

@@ Abstract
Collaborative perception has recently gained significant attention in autonomous driving, improving perception quality by enabling the exchange of additional information among vehicles. However, deploying collaborative perception systems can lead to domain shifts due to diverse environmental conditions and data heterogeneity among connected and autonomous vehicles (CAVs). To address these challenges, we propose a unified domain generalization framework to be utilized during the training and inference stages of collaborative perception. In the training phase, we introduce an Amplitude Augmentation (AmpAug) method to augment low-frequency image variations, broadening the model’s ability to learn across multiple domains. We also employ a meta-consistency training scheme to simulate domain shifts, optimizing the model with a carefully designed consistency loss to acquire domain-invariant representations. In the inference phase, we introduce an intra-system domain alignment mechanism to reduce or potentially eliminate the domain discrepancy among CAVs prior to inference. Extensive experiments substantiate the effectiveness of our method in comparison with the existing state-of-the-art works.
@@ I Introduction
Figure 1: The problem setting of domain generalization for collaborative perception, which aims to tackle the domain generalization problem in collaborative perception while aligning the domain gap among different CAVs.
@@
Recently, multi-agent collaborative perception huWhere2commCommunicationefficientCollaborative2024 ; liuWhen2comMultiAgentPerception2020 ; yangWhat2commCommunicationefficientCollaborative2023 ; chenVehicleServiceVaaS2024 ; huAgentsCoDriverLargeLanguage2024 ; xuV2XViTVehicletoEverythingCooperative2022a has attracted increasing attention in the autonomous driving community, due to its promising way to overcome the limitations of a single-agent perception system. For example, a single-agent perception system may suffer from occlusion and sparse sensor observation from afar, which may lead to a high risk of traffic accidents. Collaborative perception can address these issues by employing vehicle-to-vehicle (V2V) collaboration, where visual information (e.g., raw sensory information, intermediate perception features, and perception results) from multiple nearby CAVs can be shared to improve the accurate understanding of the environments.
@@
In collaborative perception, current approaches mainly aim to strike an optimal balance between performance and bandwidth consumption through developing new perception architectures. For example, Liu et al. liuWho2comCollaborativePerception2020 leveraged a three-step handshake communication protocol to determine the information with which CAVs should be shared. Chen et al. chenCooperCooperativePerception2019 conducted a study to implement collaborative perception by fusing raw LiDAR point clouds of different CAVs. Xu et al. xuV2XViTVehicletoEverythingCooperative2022a proposed a new vision transformer framework to fuse the information from the on-road vehicles and the roadside units (RSU). Hu et al. huAdaptiveCommunicationsCollaborative2023GLOBECOM proposed an adaptive communication scheme that can construct the communication graph and minimize the communication delay according to different channel information state (CSI).
Despite the aforementioned methods having made significant progress in balancing performance and bandwidth while optimizing communication graph construction, they regrettably overlook a critical design challenge, namely, the domain shift problem in collaborative perception. These methods typically assume that both training and testing data originate from the same domain, a condition rarely met in real-world applications.
@@
The consequence of neglecting the domain shift is a severe deterioration in perception performance when these methods are implemented in environments characterized by domain variations. For instance, in OPV2V xuOPV2VOpenBenchmark2022 and nuScenes caesarNuScenesMultimodalDataset2020 , the widely recognized benchmarks for bird’s eye view (BEV) semantic segmentation, their data collection settings exhibit notable disparities, encompassing differences in intrinsic and extrinsic parameters. One dataset is sourced from a simulator, while the other derives from real-world observations. When a model initially trained on OPV2V is directly transferred to the nuScenes dataset, a substantial performance drop is observed. Consequently, there is an urgent need to achieve domain generalization for collaborative perception.
@@
Domain generalization (DG) is an active research topic that has been investigated by many researchers before liDomainAdaptationBased2024 ; zhouDomainGeneralizationMixStyle2021 ; liDomainGeneralizationAdversarial2018 ; yangFDAFourierDomain2020 ; volpiGeneralizingUnseenDomains2018 . However, when applied to collaborative perception in the context of CAVs, the domain generalization problem becomes notably more complicated. Several challenges make collaborative perception in CAVs much more demanding for DG. First, for collaborative BEV segmentation, there are few open-source benchmarks, and annotating data is resource-intensive and time-consuming. Only OPV2V xuOPV2VOpenBenchmark2022 and V2X-Sim liV2XSimMultiAgentCollaborative2022 provide accessible datasets. Moreover, the environments of CAVs are highly dynamic, subject to frequent changes. For example, domains will shift when transitioning from city roads to country roads, and similar domain shifts will also be observed when moving from daytime to night conditions. Furthermore, data heterogeneity among CAVs also presents a significant challenge. In practical scenarios, different CAVs may operate under diverse environmental conditions during collaborative perception. For instance, one CAV might navigate in darkness, while another is positioned in a well-lit area. The absence of standardized parameters across different onboard cameras lead to variations in image quality, affecting factors like brightness, contrast, and color zhouDomainGeneralizationSurvey2022 . As a result, these divergent environmental conditions and sensor characteristics inevitably create a substantial domain gaps among CAVs, resulting in performance degradation in collaborative perception.
@@
To tackle these challenges, we present a comprehensive domain generalization framework that operates seamlessly in both training and inference phases of collaborative perception. Firstly, our approach starts with the development of an amplitude augmentation method, a resource-efficient technique that does not need extra annotated data and is ready for immediate use. Next, we introduce a meta-consistency training scheme rooted in meta-learning principle. This scheme equips the model with the ability to generalize effectively to unseen domains while fostering the acquisition of domain-invariant features. Finally, our framework includes an intra-system domain alignment scheme, strategically designed to mitigate data heterogeneity and bridge the domain gaps among various CAVs. The main contributions of this paper are summarized as follows.
@@
  • We address the practical problem of domain generalization for collaborative perception. It is the first work to study domain generalization in collaborative BEV perception.
@@
The remainder of this paper is organized as follows. In Sec. II, we review the related work of collaborative perception and domain generalization. In Sec. III, we introduce our proposed domain generalization method in detail. In Sec. IV, we conduct comprehensive experiments to evaluate our method. Finally, we conclude our work in Sec. VI.
@@ II Related Work
II-A Collaborative Perception (CP)
Despite the notable progress in autonomous driving in recent years, single-agent perception systems face significant challenges with respect to occlusions and limitations of sensor coverage. For example, in a crowded urban environment, a vehicle may be blocked by other vehicles, pedestrians, or buildings, which make it difficult to perceive its surrounding environments. Multi-agent collaborative perception (CP) emerged as a solution to addressing these issues wangV2VNetVehicletoVehicleCommunication2020a ; liuWho2comCollaborativePerception2020 ; fang2024pacp ; huCollaborativePerceptionConnected2024 .
Based on data sharing strategies, CP can generally be divided into three aspects: 1) early fusion chenCooperCooperativePerception2019 , in which raw data is shared with ego CAV; 2) intermediate fusion, in which intermediate features extracted from each CAV’s sensor data are shared; and 3) late fusion, in which perception outputs are shared (e.g., bounding box position, segmentation mask, etc.). Recent works indicate that intermediate fusion is more efficient than early fusion and late fusion because it effectively balances accuracy and communication bandwidth. For example, Liu et al. liuWho2comCollaborativePerception2020 proposed a handshake communication mechanism, which consists of three stages, namely request, match, and connect, determining which CAVs should communicate with. Li et al. liLearningDistilledCollaboration2021 leveraged knowledge distillation to train a model by a teacher-student framework and proposed a design of a matrix to allow an CAV to adaptively highlight the information region, thereby achieving effective intermediate fusion. Wang et al. wangV2VNetVehicletoVehicleCommunication2020a proposed an intermediate fusion strategy where all CAVs transmit features derived from the raw point cloud to strike a balance between bandwidth and precision. Liu et al. proposed CRCNet liuCRCNetFewshotSegmentation2022 , a few-shot segmentation network that leveraged cross-reference and local-global condition to better discover the concurrent objects in two images to enhance the few-shot segmentation learning. Wang et al. wangCollaborativeVisualNavigation2021 proposed a 3D dataset, CollaVN, to facilitate the development of multi-agent collaborative visual navigation. Moreover, Su et al. suUncertaintyQuantificationCollaborative2023 estimated the uncertainty of the object detection in CP and Hu et al. huCollaborationHelpsCamera2023 presented Coca3D to enhance camera-only 3D detection through the incorporation of multi-agent collaborations. In addition to the methods mentioned above, numerous datasets have also emerged in the field of CP to facilitate the development of this field, such as OPV2V xuOPV2VOpenBenchmark2022 , V2X-Sim liV2XSimMultiAgentCollaborative2022 , V2XSet xuV2XViTVehicletoEverythingCooperative2022a , DAIR-V2X yuDAIRV2XLargeScaleDataset2022 , and V2V4Real xuV2V4RealRealWorldLargeScale2023 . Despite the aforementioned advances of fusion methods, the domain shift caused by different environmental conditions and data heterogeneity among CAVs has not been well investigated in CP, which is the topic of this paper.
@@ II-B Domain Generalization
The domain shift problem has seriously hindered large-scale deployments of machine learning models. To tackle this issue, a lot of DG methods have been proposed gongDLOWDomainFlow2021 ; ghifaryDomainGeneralizationObject2015 ; liEpisodicTrainingDomain2019 ; yangFDAFourierDomain2020 ; liFeatureCriticNetworksHeterogeneous2019 . DG is dedicated to learning a model trained on multiple source domains that can generalize to unseen target domains. For example, Motiian et al. motiianUnifiedDeepSupervised2017 leveraged a contrastive loss to learn an embedding subspace that is discriminative, and the mapped visual domains are semantically aligned and yet maximally separated, which requires only a few labeled target samples. Some other DG methods are based on data augmentation, which is a common practice to regularize the training of machine learning models to avoid overfitting and improve generalization, e.g., Volpi et al. volpiAddressingModelVulnerability2019 defined a new data augmentation rule to transform images with different distribution where the model is most vulnerable, thereby making the model more robust against distribution shifts. Volpi et al. volpiGeneralizingUnseenDomains2018 also proposed a data augmentation method to leverage adversarial gradients obtained from a task classifier to perturb input images. In addition, meta-learning, also known as learn-to-learn, is another method helpful for DG, which exposes a model to domain shift and guides the model to learn how to learn from the domain shift. Finn et al. finnModelAgnosticMetaLearningFast2017 proposed MAML, a method that partitions the training data into meta-train and meta-test subsets, and trains a model using the meta-train subset to enhance its performance on the meta-test subset. Unfortunately, most aforementioned works are not specifically on CP. Our method also utilizes meta-learning scheme as our foundational learning scheme for CP. Here, we propose a novel amplitude augmentation method to simulate the domain shift, and leverage a consistency optimization to enable the model to learn domain-invariant features.
Figure 2: Overall Architecture. 1) The first part of our proposed method is amplitude augmentation which can transform the original source domain data to different target domains. 2) The second part is meta-consistency training, which can simulate the domain shift, guiding the model to learn how to learn from different domains. Then, we exploit the meta-consistency loss to encourage our model to learn the domain-invariant features, thereby enhancing the ability of generalization of the model. 3) The third part is the intra-system domain alignment, which can minimize the domain gap among the data perceived by different collaborative vehicles prior to inference.
@@ III-A Problem Formulation and Overview
Collaborative Domain Generalization. Let $(\mathcal{X},\mathcal{Y})$ denote the input and label space in a collaborative BEV segmentation task, $\mathcal{D}=\{\mathcal{D}_{1},\mathcal{D}_{2},\cdots,\mathcal{D}_{K}\}$ be the set of $K$ source domains. In each domain, $\mathcal{D}_{k}=\{(x_{i}^{(k)},y_{i}^{(k)})\}^{N_{k}}_{i=1}$ is a set of samples from domain distribution ($\mathcal{X}_{k},\mathcal{Y}$), where $x_{i}^{(k)}$ is the input image set of CAVs at a certain time, $y_{i}^{(k)}$ is the corresponding label, and $N_{k}$ is the number of samples in domain $k$. The goal is to learn a collaborative BEV segmentation model $F:\mathcal{X}\rightarrow\mathcal{Y}$ based on the data from multiple source domains $\mathcal{D}$, which can be generalized to the unseen target domain.
@@
Following AmpAug, we propose our meta-consistency training scheme, a method capable of simulating domain shifts and optimizing the model using a carefully crafted consistency loss function. This loss drives the encoder to learn domain-invariant features by minimizing the maximum mean discrepancy of the features obtained during both meta-training and meta-testing phases. Finally, our intra-system domain alignment mechanism is introduced, aimed at mitigating or even eliminating domain discrepancies among CAVs before the inference stage. This is achieved by translating image styles and unifying the distribution of image pixels in color space. Each of these components will be discussed in detail in the following sections. The overall framework is shown in Fig. 2.
@@ III-B Amplitude Augmentation
Figure 3: Visualization of our constructed target dataset $\mathcal{D}_{t}$ and image fast Fourier transform (FFT). Subfigure (a) gives a brief illustration of the dataset, where we observe that this dataset contains images with different weather conditions, driving scenarios, colors, etc.; Subfigure (b) visualizes FFT of images, where we transform an input image $x$ to the frequency domain by FFT and obtain the amplitude spectrum $A(x)$ (low-frequency spectrum) and phase spectrum $P(x)$ (high-frequency spectrum). The amplitude spectrum indicates the magnitude of each frequency component present in the image, which is crucial for understanding the image’s texture and style contents. The phase spectrum of an image specifies the phase or phase shifts of different spatial frequency components present in the image, providing detailed information about the spatial arrangement and positioning of features within the image.
@@
As for collaborative perception, the environments sensed by CAVs are extraordinarily complex due to dynamic transformations across time and space. Factors, such as varying weather conditions, diverse driving scenarios, fluctuating illuminations, and distinct image colors, contribute to this complexity. These environmental fluctuations result in domain discrepancy, significantly impeding perception performance when deploying models in real-world settings.
@@
Our study, upon a close examination, has revealed that this domain discrepancy primarily arises from variations in the low-frequency spectrum of images, which can be obtained by fast Fourier transform (FFT). From the perspective of image processing, the amplitude spectrum can reflect the low-frequency distribution, such as the style, which can result in the domain gap. Importantly, these low-frequency variations do not affect the image’s finer details, encompassing high-frequency spectrum elements like objects, cars, people, and textures yangFDAFourierDomain2020 ; liuFedDGFederatedDomain2021 . Furthermore, these low-frequency spectrum variations hold substantial sway over the images. If these variances are not adequately represented in the training dataset, the models fail to generalize effectively. Consequently, we introduce a straightforward yet highly effective method known as AmpAug, which serves the purpose of augmenting these low-frequency variations, compelling the model to acquire domain-invariant representations. This augmentation, in turn, reduces the model’s susceptibility to the influence of low-frequency spectrum variations in images.
@@
Based on the above observations, we derive our novel design approach. First, we construct a dataset $\mathcal{D}_{t}$, which is carefully collected from the Internet. We set some keywords to filter the images and obtain images under different weather conditions (including sunny, rainy, and foggy days), driving scenarios, colors, etc. The visualization of our constructed target dataset $\mathcal{D}_{t}$ is shown in Fig. 3. The reason for constructing this dataset is that we want to obtain a sufficient number of different amplitude spectra of the images and leverage them to augment the amplitude spectra of the images in the source domain dataset $\mathcal{D}_{s}$.
@@
Specifically, consider our constructed dataset $\mathcal{D}_{t}=\{x_{i}^{t}\}^{N_{t}}_{i=1}$ and a source domain dataset $\mathcal{D}_{s}=\{x_{i}^{s},y_{i}^{s}\}^{N_{s}}_{i=1}$, where $x_{i}^{t},x_{i}^{s}\in\mathbb{R}^{H\times W\times C}$ ($H$ and $W$ are the height and width of the image, respectively, and $C$ is the number of channels), $y_{i}^{s}\in\mathbb{R}^{H\times W\times C}$ is the corresponding label. We first sample an image $x_{i}^{t}$ from our constructed dataset $\mathcal{D}_{t}$, and convert it to the frequency domain by the fast Fourier transform (FFT):
$\mathbf{F}(x_{i}^{t})(u,v,c)=\sum^{H-1}_{h=0}\sum^{W-1}_{w=0}x_{i}^{t}(h,w,c)e^{-2j\pi\left(\frac{h}{H}u+\frac{w}{W}v\right)}.$ (1)
@@
After obtaining the frequency representation $\mathbf{F}(x_{i}^{t})$, we can decompose it into amplitude spectrum $A(x_{i}^{t})\in\mathbb{R}^{H\times W\times C}$ and phase spectrum $P(x_{i}^{t})\in\mathbb{R}^{H\times W\times C}$, which represents the low-level distributions (e.g., style, light, etc.) and high-level distributions (e.g., details and objects, etc.), respectively yangFDAFourierDomain2020 . Then, we sample another image $x_{j}^{s}$ from the source dataset and perform the same operation to obtain the amplitude spectrum $A(x_{j}^{s})$ and phase spectrum $P(x_{j}^{s})$. After that, we can obtain the augmented image $x_{j}^{s\rightarrow t}$ by incorporating the low-level distribution $A(x_{j}^{s})$ and $A(x_{i}^{t})$.
@@
Furthermore, a binary mask $M$ is introduced to control the proportion of the low-level distribution $A(x_{j}^{s})$ and $A(x_{i}^{t})$ in the augmented image $x_{j}^{s\rightarrow t}$, $M(h,w)=\mathbf{1}_{h\times w}$ where $h\in[-\alpha H:\alpha H]$, $w\in[-\alpha W:\alpha W]$, $\alpha\in(0,1)$. Then, we generate a new amplitude spectrum distribution as follows:
$A(x_{j}^{s\rightarrow t})=(I-M)\cdot A(x^{s}_{j})+M\cdot A(x^{t}_{i}),$ (2)